7. A company uses activity-based costing to determine the costs of its three products: A, B,...

7.

A company uses activity-based costing to determine the costs of its three products: A, B, and C. The budgeted cost and activity for each of the company's three activity cost pools are shown in the following table:

Budgeted Activity
Activity Cost Pool Budgeted Cost Product A Product B Product C
Activity 1 $ 70,000 6,000 9,000 20,000
Activity 2 $ 45,000 7,000 15,000 8,000
Activity 3 $ 82,000 2,500 1,000 1,625

How much overhead will be assigned to Product B using activity-based costing?

Group of answer choices

$62,500

$70,000

$197,000

$78,000

$56,500
Answer #1

Activity rate=Respective cost/Respective activity

Activity rate for:

Activity 1=70,000/(6000+9000+20,000)=$2

Activity 2=45000/(7000+15000+8000)=$1.5

Activity 3=82000/(2500+1000+1625)=$16

Hence overhead assigned to Product B=(2*9000)+(1.5*15000)+(16*1000)

=$56500